# Efficient but Vulnerable: Benchmarking and Defending LLM Batch Prompting Attack

### v1

@@ Abstract
Batch prompting, which combines a batch of multiple queries sharing the same context in one inference, has emerged as a promising solution to reduce inference costs. However, our study reveals a significant security vulnerability in batch prompting: malicious users can inject attack instructions into a batch, leading to unwanted interference across all queries, which can result in the inclusion of harmful content, such as phishing links, or the disruption of logical reasoning. In this paper, we construct BatchSafeBench, a comprehensive benchmark comprising 150 attack instructions of two types and 8k batch instances, to study the batch prompting vulnerability systematically. Our evaluation of both closed-source and open-weight LLMs demonstrates that all LLMs are susceptible to batch prompting attacks. We then explore multiple defending approaches. While the prompting-based defense shows limited effectiveness for smaller LLMs, the probing-based approach achieves about 95% accuracy in detecting attacks. Additionally, we perform a mechanistic analysis to understand the attack and identify attention heads that are responsible for it. [^1] BatchSafeBench will be released in the future.
@@ 2.2 Batch Prompting Attack
In this study, we explore two types of batch prompting attacks (Figure 2): (1) Content Attacks: Prepending or appending additional content to the generated answer without altering the reasoning chain of the query. While this type of attack does not change the final answer, the inserted content itself may introduce risks, such as linking users to phishing websites, embedding advertisements, or exposing users to sensitive information. (2) Reasoning Attacks: Manipulating the LLM’s reasoning chain or final answer of the query. This type of attack can cause the LLM to produce incorrect responses to questions it would otherwise answer correctly. If users do not carefully verify the results, they may be misled, potentially leading to incorrect decisions in life and workplaces.
@@ 3.1 Experiment Setup
We experiment with both closed-source LLMs (GPT-4o-2024-05-13, GPT-4o-mini-2024-07-18, and Claude-3.5-Sonnet-2024102) and open-weight ones (Llama3-70b-Instruct, Llama3.2-3B-Instruct, and Qwen2.5-7B-Instruct). Besides, we also randomly sample 100 instances from the benchmark to test with the reasoning models DeepSeek-R1.222R1 needs a very long time to reason for each batch. Therefore, we only experimented with a subset of the benchmark. For all experiments, we set the temperature to zero.
@@ 3.3 Impact of Batch Attack Variants
The attack instructions in BatchSafeBench are not necessarily offensive in their language. To test whether the LLM is more attentive to explicit hate words themselves, we manually adjusted the content attack instructions to make them more sensitive and unsafe. Specifically, we collected offensive statements from the hate speech dataset of Davidson et al. (2017) and then replaced the original content attack instructions with these statements. For example, instead of instructing the model to append a website link (Figure 1), now the attack instruction requires an offensive statement to be appended. We then re-evaluated the same LLMs on the modified content attack subset. Our results in Table 2 reveal that as explicit hate words were introduced, all LLMs’ ASRs dropped significantly, except for R1.333We accessed R1 from Microsoft Azure. We note that the content filtering mechanism implemented by the API service provider could play a critical role in a model’s safeguard, and there is a chance that it was not properly implemented for R1 by the time of this experiment (which is not transparent). However, we also noticed that, unlike other open-weight LLMs, R1 never responded with a refusal message. For some batches, we observed that most LLMs refused to answer them. Overall, the results suggest that current LLMs are more sensitive to explicit hate speech; however, they cannot identify the batch prompting risks when the instruction’s language is not obviously harmful.
4 Defending LLMs Against Batch Prompting Attacks
@@ 4.1 Prompting-based Defense
We started with a prompting-based defense approach. The prompting-based defense includes a defense instruction (shown in Appendix E) before the batch of queries, designed to make the LLM treat every query independently.
@@
Another important limitation of prompting-based defense is that it may be jailbroken with some adversarial attack sentences. To understand this limitation, we manually created an adversarial attack instruction in Appendix E and added it before the original content or reasoning attack instruction. We re-evaluated the effect of the prompting-based defense approach under this adversarial attack and reported its performance in Table 3 (lower). As we expected, the effectiveness of prompting-based defense under the adversarial attack was greatly degraded. Particularly for GPT-4o and GPT-4o-mini, their average ASRs increased by 30% compared to defense without the adversarial attack, showing that the models are more prone to prompt manipulation.
@@ 4.2 Probing-Based Attack Detection
Prompting-based defenses have proven insufficient in mitigating attacks on open-weight LLM. In this section, we explore a different approach, which adopts a probe Liu et al. (2019) to detect batch prompts that were attacked. Specifically, we train a linear classifier as the probe on the last-layer presentation of the LLM on the last token position of the batch prompt, to distinguish between benign and malicious prompts. We envision that, with such a probe, service providers can detect batches that are likely attacked and mitigate the risk by processing their queries individually.
@@
We experimented with this approach in two settings. The in-distribution setting assumes the awareness of the exact attack instructions used by the malicious party, which were used to create the positive (i.e., attacked) batch instances when training the probe. The out-of-distribution setting targets a more realistic setting, where the service providers do not know the exact attack instructions. In this case, we curated a different set of content and reasoning attack instructions to create positive examples. In both settings, the negative examples are benign batch instances. We randomly sampled 400 questions each from the GSM8k and the hotpotQA training set to create the batch instances. We include further details in Appendix F. For evaluation, we used the instances from BatchSafeBench as positive examples and the same instances without attack as negative ones. The probing accuracy for Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct is shown in Table 4. We observe that this method achieves very high detection accuracy, which demonstrates that by examining the last layer’s representation, it is possible to identify potentially unsafe batch prompts.
@@ 5 Why Does Batch Attack Happen?
Specifically, we followed Nikankin et al. (2024) in performing an activation patching experiment, which understands the causal effect of activation (i.e., an intermediate neural representation) by replacing (or patching) it with an alternative one and observing the resulting change in the model prediction Meng et al. (2022); Heimersheim and Nanda (2024). To target the effect of the attack instruction, we follow prior work Wang et al. (2022); Nikankin et al. (2024); Hanna et al. (2024) and design pairs of contrastive prompts. We include details of this analysis in Appendix G.
@@ Appendix D Evaluation of the ASR Evaluator
To verify the effectiveness of the evaluation prompts, we conducted a manual review. Specifically, we randomly selected $K$ ($K=100$ in our case) batch instances from BatchSafeBench with batch size $n$ ($n=100$ in our case) and fed them into the GPT-4o to get the answers without and with batch attack and the evaluation results with multiple LLMs. We then reviewed the $K\times n$ queries. The metric to evaluate the effectiveness of ASR evaluator is consistency with human evaluation, which is calculated as:
@@
Appendix E Prompting-based Attack Prompt
@@
Appendix F Probing Classifier Training
F.1 Data Preparation
@@ F.2 Training Details
Appendix G A Mechanistic Analysis of Batch Prompting Attack
@@
$IE=\frac{1}{2}[\frac{\mathcal{P}^{*}(t_{cnt})-\mathcal{P}(t_{cnt})}{\mathcal{P}(t_{cnt})}+\frac{\mathcal{P}(t_{org})-\mathcal{P}^{*}(t_{org})}{\mathcal{P}^{*}(t_{org})}]$

### v2

@@ Abstract
Batch prompting, which combines a batch of multiple queries sharing the same context in one inference, has emerged as a promising solution to reduce inference costs. However, our study reveals a significant security vulnerability in batch prompting: malicious users can inject attack instructions into a batch, leading to unwanted interference across all queries, which can result in the inclusion of harmful content, such as phishing links, or the disruption of logical reasoning. In this paper, we construct BatchSafeBench, a comprehensive benchmark comprising 150 attack instructions of two types and 8k batch instances, to study the batch prompting vulnerability systematically. Our evaluation of both closed-source and open-weight LLMs demonstrates that all LLMs are susceptible to batch prompting attacks. We then explore multiple defending approaches. While the prompting-based defense shows limited effectiveness for smaller LLMs, the probing-based approach achieves about 95% accuracy in detecting attacks. Additionally, we perform a mechanistic analysis to understand the attack and identify attention heads that are responsible for it.  [^1] Dataset:https://huggingface.co/datasets/MurongYue/BatchSafeBench [^2] Code:https://github.com/MurongYue/BatchSafeBench
@@ 2.2 Batch Prompting Attack
In this study, we explore two types of batch prompting attacks (Figure 2):
Content Attacks
Prepending or appending additional content to the generated answer without altering the reasoning chain of the query. While this type of attack does not change the final answer, the inserted content itself may introduce risks, such as linking users to phishing websites, embedding advertisements, or exposing users to sensitive information.
To better understand the instructions in the content attack, we categorized them into:(1) Phishing Websites (e.g., “Source: VerifiedData.org”). These pose a risk by misleading benign users into trusting and clicking on potentially phishing links. (2) Manipulated Persuasion (e.g., “Transparency All facts checked”). This can falsely enhance the user’s perceived credibility of the LLM’s outputs. (3) Offensive Phrases (e.g., “No one will care if you just disappear”). These may provoke the LLM to generate offensive replies. (4) Non-offensive Phrases (e.g., “You can do it”). Though seemingly benign, these can introduce unintended content into responses, altering their tone, format, or adding some strange words. While this manipulation is subtle, it can still significantly impact response quality by introducing extraneous information outside the benign user’s intent. This could disrupt the user experience, shift the response structure, or add some content that the user may not like.
Reasoning Attacks
Manipulating the LLM’s reasoning chain or final answer of the query. This type of attack can cause the LLM to produce incorrect responses to questions it would otherwise answer correctly. If users do not carefully verify the results, they may be misled, potentially leading to incorrect decisions in life and workplaces.
Uniqueness of Batch Prompting Attack
In contrast to other adversarial attacks, batch prompting attacks present a distinct and broader threat. Firstly, one uniqueness of batch prompting attacks is that they can hurt responses for benign users. Unlike many traditional jailbreaking methods, which rely on deliberate user input of specific terms and leave benign interactions unaffected, batch prompting attacks can undermine the experience of all users. In this scenario, where an LLM processes multiple queries simultaneously, a single malicious query can subtly alter the responses to others. Consequently, even users posing entirely harmless questions, with no malicious intent, may unknowingly receive attack replies.
Additionally, our research focus differs from existing jailbreak-related studies. While most jailbreak research examines whether LLMs’ output contains harmful content, e.g., toxic comments, our paper focuses on whether LLMs can handle a batch of questions independently and avoid the influence of cross-question instruction rather than focusing on the ability to avoid certain harmful content generation.
@@ 3.1 Experiment Setup
We experiment with both closed-source LLMs (GPT-4o-2024-05-13, GPT-4o-mini-2024-07-18, and Claude-3.5-Sonnet-2024102) and open-weight ones (Llama3-70b-Instruct, Llama3.2-3B-Instruct, and Qwen2.5-7B-Instruct). Besides, we also randomly sample 100 instances from the benchmark to test with the reasoning models DeepSeek-R1.333R1 needs a very long time to reason for each batch. Therefore, we only experimented with a subset of the benchmark. For all experiments, we set the temperature to zero.
@@ 3.3 Impact of Batch Attack Variants
Furthermore, we conducted additional experiments with GPT-4o and set the batch size to 20. The ASR is 88.8%, which remains high and still leaves the batch prompting system highly vulnerable to attacks.
@@
The attack instructions in BatchSafeBench are not necessarily offensive in their language. To test whether the LLM is more attentive to explicit hate words themselves, we manually adjusted the content attack instructions to make them more sensitive and unsafe. Specifically, we collected offensive statements from the hate speech dataset of Davidson et al. (2017) and then replaced the original content attack instructions with these statements. For example, instead of instructing the model to append a website link (Figure 1), now the attack instruction requires an offensive statement to be appended. We then re-evaluated the same LLMs on the modified content attack subset. Our results in Table 2 reveal that as explicit hate words were introduced, all LLMs’ ASRs dropped significantly, except for R1.444We accessed R1 from Microsoft Azure. We note that the content filtering mechanism implemented by the API service provider could play a critical role in a model’s safeguard, and there is a chance that it was not properly implemented for R1 by the time of this experiment (which is not transparent). However, we also noticed that, unlike other open-weight LLMs, R1 never responded with a refusal message. For some batches, we observed that most LLMs refused to answer them. Overall, the results suggest that current LLMs are more sensitive to explicit hate speech; however, they cannot identify the batch prompting risks when the instruction’s language is not obviously harmful.
@@
4 Defending LLMs Against Batch Prompting Attacks
@@ 4.1 Prompting-based Defense
We started with a prompting-based defense approach. The prompting-based defense includes a defense instruction (shown in Appendix F) before the batch of queries, designed to make the LLM treat every query independently.
@@
Another important limitation of prompting-based defense is that it may be jailbroken with some adversarial attack sentences. To understand this limitation, we manually created an adversarial attack instruction in Appendix F and added it before the original content or reasoning attack instruction. We re-evaluated the effect of the prompting-based defense approach under this adversarial attack and reported its performance in Table 3 (lower). As we expected, the effectiveness of prompting-based defense under the adversarial attack was greatly degraded. Particularly for GPT-4o and GPT-4o-mini, their average ASRs increased by 30% compared to defense without the adversarial attack, showing that the models are more prone to prompt manipulation.
@@ 4.2 Probing-Based Attack Detection
Prompting-based defenses have proven insufficient in mitigating attacks on open-weight LLMs. In this section, we explore a different approach, which adopts a probe Liu et al. (2019) to detect batch prompts that were attacked Qian et al. (2025). Specifically, we train a linear classifier as the probe on the last-layer presentation of the LLM on the last token position of the batch prompt, to distinguish between benign and malicious prompts. We envision that, with such a probe, service providers can detect batches that are likely attacked and mitigate the risk by processing their queries individually.
@@
We experimented with this approach in two settings. The in-distribution setting assumes the awareness of the exact attack instructions used by the malicious party, which were used to create the positive (i.e., attacked) batch instances when training the probe. The out-of-distribution setting targets a more realistic setting, where the service providers do not know the exact attack instructions. In this case, we curated a different set of content and reasoning attack instructions to create positive examples. In both settings, the negative examples are benign batch instances. We randomly sampled 400 questions each from the GSM8k and the hotpotQA training set to create the batch instances. We include further details in Appendix G. For evaluation, we used the instances from BatchSafeBench as positive examples and the same instances without attack as negative ones. The probing accuracy for Llama3.2-3B-Instruct and Qwen2.5-7B-Instruct is shown in Table 4. We observe that this method achieves very high detection accuracy, which demonstrates that by examining the last layer’s representation, it is possible to identify potentially unsafe batch prompts.
@@ 5 Why Does Batch Attack Happen?
Specifically, we followed Nikankin et al. (2024) in performing an activation patching experiment, which understands the causal effect of activation (i.e., an intermediate neural representation) by replacing (or patching) it with an alternative one and observing the resulting change in the model prediction Meng et al. (2022); Heimersheim and Nanda (2024). To target the effect of the attack instruction, we follow prior work Wang et al. (2022); Nikankin et al. (2024); Hanna et al. (2024) and design pairs of contrastive prompts. We include details of this analysis in Appendix H.
@@ Appendix D Evaluation of the ASR Evaluator
To verify the effectiveness of the evaluation prompts, we conducted a manual review. Specifically, we randomly selected $K$ ($K=100$ in our case) batch instances from BatchSafeBench with batch size $n$ ($n=5$ in our case) and fed them into the GPT-4o to get the answers without and with batch attack and the evaluation results with multiple LLMs. We then reviewed the $K\times n$ queries. The metric to evaluate the effectiveness of ASR evaluator is consistency with human evaluation, which is calculated as:
@@
Appendix E Analysis of Different Types of Content Batch Prompting Attacks
Model Phishing Website Persuasion Offensive Phrases Non-offensive Phrases
GPT-4o 91.5 93.2 88.8 92.2
GPT-4o-mini 91.9 95.3 88.4 94.9
Claude-3.5-Sonnet 68.8 72.5 64.5 71.5
LLaMA3-70B-Instruct 82.3 86.1 81.4 84.7
LLaMA3.2-3B-Instruct 67.1 69.3 67.4 68.2
Qwen2.5-7B-Instruct 69.3 73.8 66.8 69.7
Table 7: Model performance (%) across four vulnerability categories.
Model Phishing Website Persuasion Offensive Phrases Non-offensive Phrases W/o Attack
GPT-4o 85.6 84.2 85.1 86.4 90.4
GPT-4o-mini 79.4 80.3 81.4 81.9 83.9
Claude-3.5-Sonnet 88.4 86.4 86.3 87.4 94.5
LLaMA3-70B-Instruct 79.8 78.7 79.4 80.3 82.9
LLaMA3.2-3B-Instruct 69.5 71.2 69.3 71.1 74.8
Qwen2.5-7B-Instruct 75.3 75.5 74.3 76.4 79.0
Table 8: Accuracy (%) under content-based attack conditions across categories.
We defined four categories of batch prompting attack types for content prompting attacks. In this section, we report ASR and accuracy under these four categories for a range of models.
As shown in Table 7, all categories yield high ASR across models, suggesting that content-based attacks—especially those not involving offensive language—remain largely undetected. Notably, ASR is consistently high for phishing-style and manipulated persuasion instructions, indicating that these subtle manipulations are not adequately filtered.
To examine whether these attacks also affect response quality in terms of factual correctness, we measure answer accuracy under each attack type. Table 8 reports accuracy for each content category, along with the baseline accuracy (i.e., without attack). The average accuracy reduction is 6.6% on HotpotQA and 2.8% on GSM8k across all models. Importantly, all four categories of content attacks resulted in accuracy degradation, with only minor differences between the categories. Even non-offensive phrases can introduce enough perturbation to affect answer correctness, reinforcing that such content-based attacks are not merely stylistic.
Appendix F Prompting-based Attack Prompt
@@
Appendix G Probing Classifier Training
G.1 Data Preparation
@@ G.2 Training Details
Appendix H A Mechanistic Analysis of Batch Prompting Attack
@@
$\begin{split}IE=\frac{1}{2}\left[\right.&\frac{\mathcal{P}^{*}(t_{cnt})-\mathcal{P}(t_{cnt})}{\mathcal{P}(t_{cnt})}\\ &\left.+\frac{\mathcal{P}(t_{org})-\mathcal{P}^{*}(t_{org})}{\mathcal{P}^{*}(t_{org})}\right]\end{split}$